# How Does Message Passing Improve Collaborative Filtering?

### v1

@@ 3.2. Message Passing in CF Helps Low-degree Users More Compared with High-degrees
Both empirical and theoretical evidence have demonstrated that GNNs usually perform satisfactorily on high-degree nodes with rich neighbor information but not as well on low-degree nodes (tang2020investigating, ; hu2022tuneup, ). While designing graph-based model architectures for CF, most existing methods directly borrow this line of observations (wang2019neural, ; he2020lightgcn, ) and assume that the contribution of message passing for CF is similar to that for graph learning tasks in general. However, whether or not these observations still transfer to message passing in CF remains questionable, as there exist architectural and philosophical gaps between message passing for CF and its counterparts for GNNs, as discussed in Section 2. For instance, training signals for node classification usually come from labels agnostic of the graph structure; whereas the training signal for CF directly utilizes links between users and items to learn their representations. To validate these hypotheses, we conduct experiments over representative methods (i.e., LightGCN and matrix factorization (MF) trained with BPR) and show their performance w.r.t. the node degree in Figure 1.
@@
We observe that, overall both MF and LightGCN perform better on high-degree users than low-degree users. According to the upper two figures in Figure 1, MF behaves similarly to LightGCN, even without treating the input data as graphs, where the overall performance for high-degree user is stronger than that for low-degree users. However, the performance improvement of LightGCN from MF on low-degree users is larger than that for high-degree users (i.e., lower two figures in Figure 1). According to literature in general graph learning tasks (hu2022tuneup, ; liu2021tail, ; tang2020investigating, ), the performance improvement should be positively proportional to the node degree - the gain for high-degree users should be higher than that for low-degree users. This discrepancy indicates that it might not be appropriate to accredit contributions of message passing in CF directly through ideologies designed for classic graph learning tasks (e.g., node classification and link prediction).
@@
To bridge this gap, we connect supervision signals (i.e., BPR and DirectAU) commonly adopted by CF methods to Laplacian matrix learning. The formulation of BPR (rendle2009bpr, ) and DirectAU (wang2022towards, ) without the incorporation of graphs can be written as:
$\mathcal{L}_{\text{BPR}}=$ $-\sum_{(i,j)\in\mathcal{D}}\sum_{(i,k)\notin\mathcal{D}}\log\sigma(s_{ij}-s_{ik})=$
(6) $-\sum_{(i,j)\in\mathcal{D}}\sum_{(i,k)\notin\mathcal{D}}\log\sigma(\mathbf{u}_{i}^{\intercal}\cdot\mathbf{i}_{j}-\mathbf{u}_{i}^{\intercal}\cdot\mathbf{i}_{k}),$
$\mathcal{L}_{\text{DirectAU}}=$ $\sum_{(i,j)\in\mathcal{D}}||\mathbf{u}_{i}-\mathbf{i}_{j}||^{2}+\sum_{u,u^{\prime}\in\mathcal{U}}\log e^{-2||\mathbf{u}-\mathbf{u}^{\prime}||^{2}}$
(7) $+\sum_{i,i^{\prime}\in\mathcal{I}}\log e^{-2||\mathbf{i}-\mathbf{i}^{\prime}||^{2}},$
@@
where $\mathcal{D}$ refers to the set of observed interactions at the training phase and $\mathbf{i^{\prime}}$ and $\mathbf{u^{\prime}}$ refers to any random user/item. According to works on Laplacian matrix learning (zhu2021interpreting, ; dong2019learning, ; ma2021unified, ), learning node representations over graphs can be decoupled into Laplacian quadratic form, a weighted summation of two sub-goals:
(8) $\min_{\mathbf{Z}}\{||\mathbf{Z}-\mathbf{X}||^{2}+\text{tr}(\mathbf{Z}^{\intercal}\mathbf{L}\mathbf{Z})\},$
@@
where $\mathbf{Z}$ refers to the node representation matrix after the message passing, $\mathbf{X}$ refers to the input feature matrix, and $\mathbf{L}$ refers to the Laplacian matrix. The first term regularizes the latent representation such that it does not diverge too much from the input feature; whereas the second term promotes the similarity between latent representations of adjacent nodes, which can be re-written as: $\text{tr}(\mathbf{Z}^{\intercal}\cdot\mathbf{L}\cdot\mathbf{Z})=\sum_{(i,j)\in\mathcal{D}}||\mathbf{u}_{i}-\mathbf{i}_{j}||^{2}$ in CF bipartite graphs. (zhu2021interpreting, ) show that $K$ layers of linear message passing exactly optimizes the second term in Equation 8. Given this theoretical foundation, we derive the following theorem w.r.t. relations between BPR, DirectAU, and message passing in CF:
Theorem 1 ().
[…]
@@
Combining this theoretical relation with the aforementioned empirical observations, we show that these two supervision signals could inadvertently conduct message passing in the backward step, even without explicitly treating interaction data as graphs. Since this inadvertent message passing happens during the back-propagation, its performance is positively correlated to the amount of training signals a user/item can get. In the case of CF, the amount of training signals for a user is directly proportional to the node degree. High-degree active users naturally benefit more from the inadvertent message passing from objective functions, because they acquire more training signals from the objective function. Hence, when explicit message passing is applied to CF methods, the performance gain for high-degree users is less significant than that for low-degree users. Because the contribution of the message passing over high-degree nodes has been mostly fulfilled by the inadvertent message passing during the training.
To quantitatively prove this theory, we incrementally upsample low-degree training users and observe the performance improvement that TAG-CF could introduce at each upsampling rate. If our line of theory is correct, then we should expect less performance improvement on low-degree users for a larger upsampling rate. The results are shown in Section 5.6 with supporting evidence.
4. Test-time Aggregation for Collaborative Filtering
@@
The test-time aggregation is inspired by our first perspective that, within total performance gains brought by message passing, gains from additional neighbor representations during the forward pass dominate those brought by accompanying gradient updates to neighbors during the back-propagation. Applying message passing only at test time avoids repetitive training-time queries (i.e., once per node and epoch) of surrounding neighbors, which grow exponentially as the number of layers increases by the neighbor explosion phenomenon (guo2023linkless, ; zhang2021graph, ; zeng2021decoupling, ). Specifically, given a set of well-trained user and item representations $\mathbf{U}\in\mathbb{R}^{|\mathcal{U}|\times d}$ and $\mathbf{I}\in\mathbb{R}^{|\mathcal{I}|\times d}$, TAG-CF augments representations for user $u_{i}$ and item $i_{i}$ as:
(9) $\mathbf{u}_{i}^{*}=\mathbf{u}_{i}+\sum_{i_{j}\in N(u_{i})}|N(u_{i})|^{m}|N(i_{j})|^{n}\cdot\mathbf{i}_{j},$
$\mathbf{i}_{i}^{*}=\mathbf{i}_{i}+\sum_{u_{j}\in N(i_{i})}|N(i_{i})|^{m}|N(u_{j})|^{n}\cdot\mathbf{u}_{j},$
@@
where $m$ and $n$ are two hyper-parameters that control the normalization of message passing. With $m=n=-\frac{1}{2}$, Equation 9 becomes the exact formulation of one-layer LightGCN (i.e., Equation 3). Empirically, we observe that the setup with $m=n=-\frac{1}{2}$ for TAG-CF does not always work for all datasets. This setup is directly migrated from message passing for homogeneous graphs (kipf2016semi, ), which might not be applicable for bipartite graphs where all neighbors are heterogeneous (dasoulas2020learning, ). Unlike LightGCN which can fill this gap by adaptively tuning all representations during the training, TAG-CF cannot update any parameter since it is applied at test time, and hence requires tune-able normalization hyper-parameters.
@@
Moreover, following our second perspective that message passing helps low-degree nodes more in CF, we further derive ${}^{+}$, which reduces the cost of TAG-CF by applying the one-time message passing only to low-degree nodes with sparse neighborhoods. Focusing on only low-degree nodes has two benefits: (i) it reduces the number of nodes that ${}^{+}$ needs to attend to, and (ii) message passing for low-degree nodes is naturally cheaper than for high-degree nodes given the surrounding neighborhoods are sparser (mitigating neighbor explosion). The degree threshold that determines which nodes to apply ${}^{+}$ is selected by the validation performance, with details in Appendix C.
@@
Leveraging our novel takeaways, TAG-CF can effectively enhance MF methods by conducting message passing only once at test time. The key novelty of TAG-CF is that it effectively utilizes graphs while circumventing most of notorious computational overheads of message passing. It is extremely flexible, simple to implement, and enjoys the performance benefits of graph-based CF method while paying the lowest overall scalability.
5. Experiments
@@
We conduct extensive experiments to demonstrate the effectiveness and efficiency of TAG-CF. Specifically, we aim to answer the following research questions: RQ (1): how effective is TAG-CF at improving MF methods without using graphs, RQ (2): how much overheads does TAG-CF introduce, RQ (3): can TAG-CF effectively enhance MF methods trained by different objectives, RQ (4): how effective is ${}^{+}$ w.r.t. different degree cutoffs, and RQ (5): do behaviors of TAG-CF align with our findings in Section 3?
@@
Table 2. Recommendation performance (i.e., NDCG@20 and Recall@20) of all models across users with different numbers of interactions. The lower percentile indicates the set of nodes whose degrees are ranked in the lower 30% population. Bold and underline indicate the best and second best model respectively. LightGCN and MF are trained with DirectAU (wang2022towards, ).
@@
Amazon-Book 5.32±${}_{\pm{0.08}}$ 8.09±${}_{\pm{0.10}}$ 5.33±${}_{\pm{0.02}}$ 5.67±${}_{\pm{0.03}}$ 6.4% 8.02±${}_{\pm{0.07}}$ 8.26±${}_{\pm{0.06}}$ 3.0% 5.61±${}_{\pm{0.19}}$ 6.04±${}_{\pm{0.21}}$ 7.7%
Anime 20.13±${}_{\pm{0.18}}$ 27.78±${}_{\pm{0.21}}$ 22.23±${}_{\pm{0.19}}$ 22.58±${}_{\pm{0.15}}$ 1.6% 23.95±${}_{\pm{0.07}}$ 27.15±${}_{\pm{0.04}}$ 13.4% 28.14±${}_{\pm{0.19}}$ 30.10±${}_{\pm{0.21}}$ 7.0%
Gowalla 8.46±${}_{\pm{0.06}}$ 10.08±${}_{\pm{0.13}}$ 3.87±${}_{\pm{0.15}}$ 4.08±${}_{\pm{0.11}}$ 5.4% 10.00±${}_{\pm{0.08}}$ 10.19±${}_{\pm{0.04}}$ 1.9% 8.21±${}_{\pm{0.09}}$ 8.63±${}_{\pm{0.11}}$ 5.1%
Yelp-2018 4.87±${}_{\pm{0.06}}$ 6.10±${}_{\pm{0.09}}$ 3.11±${}_{\pm{0.07}}$ 3.26±${}_{\pm{0.04}}$ 4.8% 6.08±${}_{\pm{0.08}}$ 6.18±${}_{\pm{0.05}}$ 1.7% 4.89±${}_{\pm{0.10}}$ 5.44±${}_{\pm{0.12}}$ 11.2%
MovieLens-1M 22.13±${}_{\pm{0.26}}$ 25.95±${}_{\pm{0.28}}$ 18.34±${}_{\pm{0.19}}$ 22.53±${}_{\pm{0.21}}$ 22.8% 20.98±${}_{\pm{0.12}}$ 29.20±${}_{\pm{0.19}}$ 39.2% 23.89±${}_{\pm{0.19}}$ 28.37±${}_{\pm{0.21}}$ 18.8%
Internal 5.91±${}_{\pm{0.07}}$ 8.12±${}_{\pm{0.03}}$ OOM - - 6.79±${}_{\pm{0.04}}$ 8.52±${}_{\pm{0.06}}$ 25.5% OOM - -
@@
Amazon-Book 6.97±${}_{\pm{0.11}}$ 8.06±${}_{\pm{0.11}}$ 6.13±${}_{\pm{0.13}}$ 6.54±${}_{\pm{0.09}}$ 6.7% 8.01±${}_{\pm{0.03}}$ 8.13±${}_{\pm{0.03}}$ 1.5% 5.77±${}_{\pm{0.25}}$ 6.11±${}_{\pm{0.27}}$ 5.9%
Anime 22.54±${}_{\pm{0.25}}$ 27.97±${}_{\pm{0.21}}$ 30.17±${}_{\pm{0.09}}$ 30.86±${}_{\pm{0.12}}$ 2.3% 24.01±${}_{\pm{0.06}}$ 27.25±${}_{\pm{0.03}}$ 9.8% 30.30±${}_{\pm{0.11}}$ 30.89±${}_{\pm{0.11}}$ 1.9%
Gowalla 8.65±${}_{\pm{0.10}}$ 9.96±${}_{\pm{0.11}}$ 5.23±${}_{\pm{0.04}}$ 5.29±${}_{\pm{0.05}}$ 1.1% 9.77±${}_{\pm{0.08}}$ 9.88±${}_{\pm{0.04}}$ 1.1% 8.53±${}_{\pm{0.14}}$ 9.02±${}_{\pm{0.15}}$ 5.7%
Yelp-2018 5.54±${}_{\pm{0.06}}$ 6.33±${}_{\pm{0.06}}$ 3.79±${}_{\pm{0.09}}$ 3.89±${}_{\pm{0.05}}$ 2.6% 6.25±${}_{\pm{0.06}}$ 6.36±${}_{\pm{0.03}}$ 1.8% 5.01±${}_{\pm{0.11}}$ 5.53±${}_{\pm{0.11}}$ 10.4%
MovieLens-1M 23.17±${}_{\pm{0.18}}$ 26.64±${}_{\pm{0.23}}$ 20.57±${}_{\pm{0.18}}$ 22.98±${}_{\pm{0.20}}$ 11.7% 22.51±${}_{\pm{0.14}}$ 29.65±${}_{\pm{0.17}}$ 31.7% 26.50±${}_{\pm{0.15}}$ 29.68±${}_{\pm{0.21}}$ 12.0%
Internal 6.94±${}_{\pm{0.06}}$ 8.10±${}_{\pm{0.06}}$ OOM - - 7.04±${}_{\pm{0.02}}$ 8.54±${}_{\pm{0.02}}$ 21.3% OOM - -
@@
Amazon-Book 10.71±${}_{\pm{0.14}}$ 13.18±${}_{\pm{0.17}}$ 10.42±${}_{\pm{0.16}}$ ${}_{\pm{0.11}}$ 6.3% 13.07±${}_{\pm{0.09}}$ 13.37±${}_{\pm{0.10}}$ 2.3% 7.92±${}_{\pm{0.15}}$ 8.31±${}_{\pm{0.10}}$ 4.9%
Anime 25.74±${}_{\pm{0.35}}$ 32.74±${}_{\pm{0.21}}$ 37.14±${}_{\pm{0.59}}$ 38.41±${}_{\pm{0.53}}$ 3.4% 29.08±${}_{\pm{0.09}}$ 31.94±${}_{\pm{0.05}}$ 9.8% 33.96±${}_{\pm{0.28}}$ 36.49±${}_{\pm{0.28}}$ 7.4%
Gowalla 17.53±${}_{\pm{0.32}}$ 19.14±${}_{\pm{0.20}}$ 8.73±${}_{\pm{0.08}}$ 9.01±${}_{\pm{0.06}}$ 3.2% 18.92±${}_{\pm{0.19}}$ 19.17±${}_{\pm{0.13}}$ 1.3% 15.57±${}_{\pm{0.18}}$ 16.01±${}_{\pm{0.15}}$ 2.8%
Yelp-2018 10.15±${}_{\pm{0.13}}$ 10.75±${}_{\pm{0.14}}$ 7.17±${}_{\pm{0.06}}$ 7.54±${}_{\pm{0.12}}$ 5.2% 10.63±${}_{\pm{0.13}}$ 10.98±${}_{\pm{0.14}}$ 3.3% 7.71±${}_{\pm{0.15}}$ 8.59±${}_{\pm{0.18}}$ 11.4%
MovieLens-1M 22.71±${}_{\pm{0.16}}$ 25.80±${}_{\pm{0.22}}$ 19.58±${}_{\pm{0.14}}$ 24.11±${}_{\pm{0.16}}$ 23.1% 23.64±${}_{\pm{0.18}}$ 28.10±${}_{\pm{0.20}}$ 18.9% 26.13±${}_{\pm{0.21}}$ 28.97±${}_{\pm{0.23}}$ 10.9%
Internal 10.54±${}_{\pm{0.09}}$ 13.81±${}_{\pm{0.02}}$ OOM - - 11.13±${}_{\pm{0.05}}$ 13.97±${}_{\pm{0.06}}$ 25.5% OOM - -

### v2

@@ 3.2 Message Passing in CF Helps Low-degree Users More Compared with High-degrees
Both empirical and theoretical evidence have demonstrated that GNNs usually perform satisfactorily on high-degree nodes with rich neighbor information but not as well on low-degree nodes [54, 23]. While designing graph-based model architectures for CF, most existing methods directly borrow this line of observations [60, 20] and assume that the contribution of message passing for CF is similar to that for graph learning tasks in general. However, whether or not these observations still transfer to message passing in CF remains questionable, as there exist architectural and philosophical gaps between message passing for CF and its counterparts for GNNs, as discussed in Section 2. To validate these hypotheses, we conduct experiments over representative methods (i.e., LightGCN and matrix factorization (MF) trained with BPR) and show their performance w.r.t. the node degree in Figure 1.
@@
We observe that, overall both MF and LightGCN perform better on high-degree users than low-degree users. According to the upper two figures in Figure 1, MF behaves similarly to LightGCN, even without treating the input data as graphs, where the overall performance for high-degree user is stronger than that for low-degree users. However, the performance improvement of LightGCN from MF on low-degree users is larger than that for high-degree users (i.e., lower two figures in Figure 1). According to literature in general graph learning tasks [23, 36, 54], the performance improvement should be positively proportional to the node degree - the gain for high-degree users should be higher than that for low-degree users. This discrepancy indicates that it might not be appropriate to accredit contributions of message passing in CF directly through ideologies designed for classic graph learning tasks (e.g., node classification and link prediction).
Figure 1: Performances of LightGCN and Matrix Factorization w.r.t. the user degree across datasets. The performance improvement brought by message passing decreases as the user degree goes up.
@@
To bridge this gap, we connect supervision signals (i.e., BPR and DirectAU) commonly adopted by CF methods to Laplacian matrix learning. The formulation of BPR [45] and DirectAU [57] without the incorporation of graphs can be written as:
$\mathcal{L}_{\text{BPR}}=-\sum_{(i,j)\in\mathcal{D}}\sum_{(i,k)\notin\mathcal{D}}\log\sigma(s_{ij}-s_{ik})=-\sum_{(i,j)\in\mathcal{D}}\sum_{(i,k)\notin\mathcal{D}}\log\sigma(\mathbf{u}_{i}^{\intercal}\cdot\mathbf{i}_{j}-\mathbf{u}_{i}^{\intercal}\cdot\mathbf{i}_{k}),$
$\mathcal{L}_{\text{DirectAU}}=\sum_{(i,j)\in\mathcal{D}}||\mathbf{u}_{i}-\mathbf{i}_{j}||^{2}+\sum_{u,u^{\prime}\in\mathcal{U}}\log e^{-2||\mathbf{u}-\mathbf{u}^{\prime}||^{2}}+\sum_{i,i^{\prime}\in\mathcal{I}}\log e^{-2||\mathbf{i}-\mathbf{i}^{\prime}||^{2}},$ (5)
@@
where $\mathcal{D}$ refers to the set of observed interactions at the training phase and $\mathbf{i^{\prime}}$ and $\mathbf{u^{\prime}}$ refers to any random user/item. According to works on Laplacian matrix learning [82, 10, 38], learning node representations over graphs can be decoupled into Laplacian quadratic form, a weighted summation of two sub-goals:
$\min_{\mathbf{Z}}\{||\mathbf{Z}-\mathbf{X}||^{2}+\text{tr}(\mathbf{Z}^{\intercal}\mathbf{L}\mathbf{Z})\},$ (6)
@@
where $\mathbf{Z}$ refers to the node representation matrix after the message passing, $\mathbf{X}$ refers to the input feature matrix, and $\mathbf{L}$ refers to the Laplacian matrix. The first term regularizes the latent representation such that it does not diverge too much from the input feature; whereas the second term promotes the similarity between latent representations of adjacent nodes, which can be re-written as: $\text{tr}(\mathbf{Z}^{\intercal}\cdot\mathbf{L}\cdot\mathbf{Z})=\sum_{(i,j)\in\mathcal{D}}||\mathbf{u}_{i}-\mathbf{i}_{j}||^{2}$ in CF bipartite graphs. [82] show that $K$ layers of linear message passing exactly optimizes the second term in Equation 6. Given this theoretical foundation, we derive the following theorem w.r.t. relations between BPR, DirectAU, and message passing in CF:
Theorem 1.
Assuming that $||\mathbf{u}_{i}||^{2}=||\mathbf{i}_{j}||^{2}=1$ for any $u_{i}\in\mathcal{U}$ and $I_{j}\in\mathcal{I}$, objectives of BPR and DirectAU are strictly upper-bounded by the objective of message passing (i.e., $\mathcal{L}_{\text{BPR}}\leq\sum_{(i,j)\in\mathcal{D}}||\mathbf{u}_{i}-\mathbf{i}_{j}||^{2}$ and $\mathcal{L}_{\text{DirectAU}}\leq\sum_{(i,j)\in\mathcal{D}}||\mathbf{u}_{i}-\mathbf{i}_{j}||^{2}$).
@@
Combining this theory with the aforementioned empirical observations, we show that these two supervision signals could inadvertently conduct message passing in the backward step, even without explicitly treating interaction data as graphs. Since this inadvertent message passing happens during the back-propagation, its performance is positively correlated to the amount of training signals a user/item can get. In the case of CF, the amount of training signals for a user is directly proportional to the node degree. High-degree active users naturally benefit more from the inadvertent message passing from objective functions, because they acquire more training signals. Hence, when explicit message passing is applied to CF methods, the performance gain for high-degree users is less significant than that for low-degree users. Because the contribution of the message passing over high-degree nodes has been mostly fulfilled by the inadvertent message passing during the training.
To quantitatively prove this theory, we incrementally upsample low-degree training users and observe the performance improvement that TAG-CF could introduce at each upsampling rate. If our line of theory is correct, then we should expect less performance improvement on low-degree users for a larger upsampling rate. The results are shown in Appendix E with supporting evidence.
4 Test-time Aggregation for Collaborative Filtering
@@
The test-time aggregation is inspired by our first perspective that, within total performance gains brought by message passing, gains from additional neighbor representations during the forward pass dominate those brought by accompanying gradient updates to neighbors during the back-propagation. Applying message passing only at test time avoids repetitive training-time queries (i.e., once per node and epoch) of surrounding neighbors, which grow exponentially as the number of layers increases by the neighbor explosion phenomenon [16, 76, 75]. Specifically, given a set of well-trained user and item representations $\mathbf{U}\in\mathbb{R}^{|\mathcal{U}|\times d}$ and $\mathbf{I}\in\mathbb{R}^{|\mathcal{I}|\times d}$, TAG-CF augments representations for user $u_{i}$ and item $i_{i}$ as:
$\mathbf{u}_{i}^{*}=\mathbf{u}_{i}+\sum_{i_{j}\in N(u_{i})}|N(u_{i})|^{m}|N(i_{j})|^{n}\cdot\mathbf{i}_{j},\mathbf{i}_{i}^{*}=\mathbf{i}_{i}+\sum_{u_{j}\in N(i_{i})}|N(i_{i})|^{m}|N(u_{j})|^{n}\cdot\mathbf{u}_{j},$ (7)
@@
where $m$ and $n$ are two hyper-parameters that control the normalization of message passing. With $m=n=-\frac{1}{2}$, Equation 7 becomes the exact formulation of one-layer LightGCN (i.e., Equation 2). Empirically, we observe that the setup with $m=n=-\frac{1}{2}$ for TAG-CF does not always work for all datasets. This setup is directly migrated from message passing for homogeneous graphs [30], which might not be applicable for bipartite graphs where all neighbors are heterogeneous [7]. Unlike LightGCN which can fill this gap by adaptively tuning all representations during the training, TAG-CF cannot update any parameter since it is applied at test time, and hence requires tune-able normalization hyper-parameters.
@@
Moreover, following our second perspective that message passing helps low-degree nodes more in CF, we further derive TAG-CF+, which reduces the cost of TAG-CF by applying the one-time message passing only to low-degree nodes. Focusing on only low-degree nodes has two benefits: (i) it reduces the number of nodes that TAG-CF+ needs to attend to, and (ii) message passing for low-degree nodes is naturally cheaper than for high-degree nodes given the surrounding neighborhoods are sparser (mitigating neighbor explosion). The degree threshold that determines which nodes to apply TAG-CF+ is selected by the validation performance, with details in Appendix C.
@@
TAG-CF can effectively enhance MF methods by conducting message passing only once at test time. TAG-CF effectively utilizes graphs while circumventing most of notorious computational overheads of message passing. It is extremely flexible, simple to implement, and enjoys the performance benefits of graph-based CF method while paying the lowest overall scalability.
5 Experiments
@@
We conduct extensive experiments to demonstrate the effectiveness and efficiency of TAG-CF. We aim to answer the following research questions: RQ (1): how effective is TAG-CF at improving MF methods without using graphs, RQ (2): how much overheads does TAG-CF introduce, RQ (3): can TAG-CF effectively enhance MF methods trained by different objectives, RQ (4): how effective is TAG-CF+ w.r.t. different degree cutoffs, and RQ (5): do behaviors of TAG-CF align with our findings in Section 3?
5.1 Experimental Settings
Baselines. We compare TAG-CF with two branches of methods: (1) CF methods that do not explicitly utilize graphs, including vanilla matrix factorization (MF) methods trained by BPR and DirectAU [45, 57], Efficient Neural Matrix Factorization [3] (denoted as ENMF), and UltraGCN [39]. (2) Graph-based CF methods, including LightGCN [20] and NGCF [60]. Besides, we also compare with recent graph-based CF methods that extend LightGCN by adding additional self-supervised signals for better performance, including LightGCL [1], SimGCL [74], and SGL [65]. For the coherence of reading, we include comprehensive discussions about evaluation protocols across all methods, tuning for hyper-parameters, and other implementation details in Appendix C.
5.2 Performance Improvement to Matrix Factorization Methods
@@
Table 2: Recommendation performance (i.e., NDCG@20 and Recall@20) of all models across users with different numbers of interactions. The lower percentile indicates the set of nodes whose degrees are ranked in the lower 30% population. Bold and underline indicate the best and second best model respectively. LightGCN and MF are trained with DirectAU [57].
@@
Amazon-Book 5.32±0.08 8.09±0.10 5.33±0.02 5.67±0.03 6.4% 8.02±0.07 8.26±0.06 3.0% 5.61±0.19 6.04±0.21 7.7%
Anime 20.13±0.18 27.78±0.21 22.23±0.19 22.58±0.15 1.6% 23.95±0.07 27.15±0.04 13.4% 28.14±0.19 30.10±0.21 7.0%
Gowalla 8.46±0.06 10.08±0.13 3.87±0.15 4.08±0.11 5.4% 10.00±0.08 10.19±0.04 1.9% 8.21±0.09 8.63±0.11 5.1%
Yelp-2018 4.87±0.06 6.10±0.09 3.11±0.07 3.26±0.04 4.8% 6.08±0.08 6.18±0.05 1.7% 4.89±0.10 5.44±0.12 11.2%
MovieLens-1M 22.13±0.26 25.95±0.28 18.34±0.19 22.53±0.21 22.8% 20.98±0.12 29.20±0.19 39.2% 23.89±0.19 28.37±0.21 18.8%
Internal 5.91±0.07 8.12±0.03 OOM - - 6.79±0.04 8.52±0.06 25.5% OOM - -
@@
Amazon-Book 6.97±0.11 8.06±0.11 6.13±0.13 6.54±0.09 6.7% 8.01±0.03 8.13±0.03 1.5% 5.77±0.25 6.11±0.27 5.9%
Anime 22.54±0.25 27.97±0.21 30.17±0.09 30.86±0.12 2.3% 24.01±0.06 27.25±0.03 9.8% 30.30±0.11 30.89±0.11 1.9%
Gowalla 8.65±0.10 9.96±0.11 5.23±0.04 5.29±0.05 1.1% 9.77±0.08 9.88±0.04 1.1% 8.53±0.14 9.02±0.15 5.7%
Yelp-2018 5.54±0.06 6.33±0.06 3.79±0.09 3.89±0.05 2.6% 6.25±0.06 6.36±0.03 1.8% 5.01±0.11 5.53±0.11 10.4%
MovieLens-1M 23.17±0.18 26.64±0.23 20.57±0.18 22.98±0.20 11.7% 22.51±0.14 29.65±0.17 31.7% 26.50±0.15 29.68±0.21 12.0%
Internal 6.94±0.06 8.10±0.06 OOM - - 7.04±0.02 8.54±0.02 21.3% OOM - -
@@
Amazon-Book 10.71±0.14 13.18±0.17 10.42±0.16 11.08±0.11 6.3% 13.07±0.09 13.37±0.10 2.3% 7.92±0.15 8.31±0.10 4.9%
Anime 25.74±0.35 32.74±0.21 37.14±0.59 38.41±0.53 3.4% 29.08±0.09 31.94±0.05 9.8% 33.96±0.28 36.49±0.28 7.4%
Gowalla 17.53±0.32 19.14±0.20 8.73±0.08 9.01±0.06 3.2% 18.92±0.19 19.17±0.13 1.3% 15.57±0.18 16.01±0.15 2.8%
Yelp-2018 10.15±0.13 10.75±0.14 7.17±0.06 7.54±0.12 5.2% 10.63±0.13 10.98±0.14 3.3% 7.71±0.15 8.59±0.18 11.4%
MovieLens-1M 22.71±0.16 25.80±0.22 19.58±0.14 24.11±0.16 23.1% 23.64±0.18 28.10±0.20 18.9% 26.13±0.21 28.97±0.23 10.9%
Internal 10.54±0.09 13.81±0.02 OOM - - 11.13±0.05 13.97±0.06 25.5% OOM - -